# Rémi A. Chou and Matthieu R. Bloch

### v1

@@ VI-A Achievability
Due to the similarity with the achievability proof of Theorem 1, we only highlight the main steps of the proof.
@@ VI-A2 Concealment
$\mathbb{V}(q_{KF\bar{Y}_{b}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}q_{\bar{Y}_{b}^{n}})$ $\leqslant 2\epsilon+\sqrt{2^{r-\bar{n}H(X|Y_{b})+\bar{n}\delta(\bar{n})}}$
@@
From (59) and [20, Lemma 2.7], we conclude that the concealment requirement (38) holds with $r$ such that $\lim_{n\to\infty}\frac{r}{n}\leqslant\min_{b\in\mathcal{B}}H(X|Y_{b})-\delta$, $\delta>0$.
@@ VI-B Converse
Due to the similarity with the converse proof of Theorem 2, we again only highlight the main steps of the proof.
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}I(AS;Y_{\mathcal{B}}^{n}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}\bar{M}_{n,1:r_{n}-1}M_{n,r_{n}}X^{n})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}I(AS;Y_{\mathcal{B}}^{n}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}\bar{M}_{n,1:r_{n}-1}X^{n})$ (61)
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}I(AS;Y_{\mathcal{B}}^{n-1}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}X^{n})$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}I(AS;Y_{\mathcal{B}}^{n-1}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}X^{n-1})$ (62)
@@
    holds by repeating $n-1$ times the steps between (60) and (62).
@@ VII Concluding Remarks
We investigated multi-user commitment in two distinct settings. In the first setting, a verifier interacts with $L$ bidders, each committing to individual messages. This setting explores whether a multi-bidder protocol can outperform single-bidder protocols and time-sharing. Our results answer this question positively, and establish the sum-rate capacity when the bidders are colluding and the capacity region when they are non-colluding. In the second setting, a single bidder can interact with multiple verifiers. Our results characterize the commitment capacity, showing that positive commitment rates can still be achieved even if some verifiers drop out of the network after the commit phase, a scenario where commitment would otherwise be impossible with only a single verifier.

### v2

@@ VI-A Achievability
Because of the similarity with the achievability proof of Theorem 1, we only highlight the main steps of the proof.
@@ VI-A2 Concealment
$\mathbb{V}(q_{KF\bar{Y}_{b}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}q_{\bar{Y}_{b}^{n}})$
$\leqslant 2\epsilon+\sqrt{2^{r-\bar{n}H(X|Y_{b})+\bar{n}\delta(\bar{n})}}$
@@
From (61) and Lemma 3, we conclude that the concealment requirement (38) holds with $r$ such that $\lim_{n\to\infty}\frac{r}{n}\leqslant\min_{b\in\mathcal{B}}H(X|Y_{b})-\delta$, $\delta>0$.
@@ VI-B Converse
Again, because of the similarity with the converse proof of Theorem 2, we again only highlight the main steps of the proof.
@@
$\stackrel{{\scriptstyle(a)}}{{=}}I(AS;Y_{\mathcal{B}}^{n}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}\bar{M}_{n,1:r_{n}-1}M_{n,r_{n}}X^{n})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}I(AS;Y_{\mathcal{B}}^{n}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}\bar{M}_{n,1:r_{n}-1}X^{n})$ (63)
@@
$\stackrel{{\scriptstyle(d)}}{{=}}I(AS;Y_{\mathcal{B}}^{n-1}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}X^{n})$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}I(AS;Y_{\mathcal{B}}^{n-1}S^{\prime}_{\mathcal{B}}|\bar{M}_{1:n-1}X^{n-1})$ (64)
@@
    holds by repeating n1n-1 times the steps between (62) and (64).
@@ VII Concluding Remarks
We have investigated multi-user commitment in two distinct settings. In the first setting, a verifier interacts with $L$ bidders, each committing to individual messages. This setting explores whether a multi-bidder protocol can outperform single-bidder protocols and time-sharing. Our results answer this question positively, and establish the sum-rate capacity when the bidders are colluding and the capacity region when they are non-colluding. In the second setting, a single bidder can interact with multiple verifiers. Our results characterize the commitment capacity, showing that positive commitment rates can still be achieved even if some verifiers drop out of the network after the commit phase, a scenario where commitment would otherwise be impossible with only a single verifier.